Eleven years ago, Max was three times as old as his son. Three years from now, he will be twice as old as his son. Find the current ages of Max and his son.
Answer by Maths68(1474)

Let
Current age of Max=m
Current age of his son=s
Eleven years ago means we have to subtract 11 from their current ages
m-11=s-11
Then
Max was three times as old as his son
m-11=3(s-11)
m-11=3s-33
m=3s-33+11
m=3s-22 ..........(1)
Three years from now means we have to add 3 to their current ages
m+3=s+3
Then Max will be twice as old as his son
m+3=2(s+3)
m+3=2s+6
m=2s+6-3
m=2s+3..........(2)
Put the value of m from (1) to (2)
3s-22=2s+3
3s-2s=3+22
s=25
Put the value of s in (1) or (2)
m=2(25)+3..........(2)
m=50+3
m=53
Max is 53 and his son is 25